OpenStudy (samigupta8):

a+b=3-cos 4θ a-b=4 sin 2θ then ab is always less than or equal to p.find the value of p

OpenStudy (loser66):

You do I guide, deal?

OpenStudy (samigupta8):

kkk....

OpenStudy (loser66):

It's my order, not options

OpenStudy (samigupta8):

okkk.....

OpenStudy (loser66):

plus 2 equations together and solve for a minus 2 equations together and solve for b time a *b set it =p , done

OpenStudy (loser66):

* set it less than or equal to p

OpenStudy (loser66):

My job is done, now, you perform. YOU CAN DO IT.

OpenStudy (samigupta8):

it's cumin as a=3-cos4θ+4 sin 2θ/2 and b=3-cos 4θ-4 sin 2θ/2

OpenStudy (loser66):

put parentheses in, to not confuse just sin 2theta/2. got what I mean? and then, GO AHEAD

OpenStudy (samigupta8):

i didn't get what u wanted 2 say

OpenStudy (loser66):

\[a= \frac{3-cos4\theta + 4sin 2\theta}{2}\]the whole thing over 2. not just the last one. that's what I mean

OpenStudy (samigupta8):

YES..YYESS'

OpenStudy (samigupta8):

THEN WHAT I HAVE TO DO

OpenStudy (loser66):

a*b that's it

OpenStudy (samigupta8):

u mean that i have to multiply each and evry term of it and that will give me the ans

OpenStudy (loser66):

when you time a*b , the condition tells you that ab<=p so, you time them, and put them into the form ab<=p. that's it. or p >= ab, that's the same/

OpenStudy (loser66):

@dan815 check mine, dan

OpenStudy (dan815):